How many grams of FeCl3 that will form if 1.0 g of each reactant mix?


1
Expert's answer
2021-02-22T06:05:52-0500

2 Fe + 3 Cl2 "\to" 2 FeCl3

Moles of Fe = 1g/56g•mol-1 = 0.0178 moles

2 moles of Fe = 2 moles of FeCl3

Grams of FeCl3 = 0.0178× 162.2 = 2.88g